An investment project provides cash inflows of $650 per year for 11 years.
a. What is the project payback period if the initial cost is $2,600?

  

b. What is the project payback period if the initial cost is $4,290?

  

c. What is the project payback period if the initial cost is $7,800?
Answer #1

payback period= initial cost/annual cash flows

1.payback period=2,600/650=4 years

2.payback period=4290/650=6.6 years

3.payback period=7800/650=Never.